### v1

@@
Abstract
Fast and accurate depth sensing has long been a significant research challenge. Event camera, as a device that quickly responds to intensity changes, provides a new solution for structured light (SL) systems. In this paper, we introduce Gray code into event-based SL systems for the first time. Our setup includes an event camera and Digital Light Processing (DLP) projector, enabling depth estimation through high-speed projection and decoding of Gray code patterns. By employing spatio-temporal encoding for point matching, our method is immune to timestamp noise, realizing high-speed depth estimation without loss of accuracy. The binary nature of events and Gray code minimizes data redundancy, enabling us to fully utilize sensor bandwidth at 100%. Experimental results show that our approach achieves accuracy comparable to state-of-the-art scanning methods while surpassing them in data acquisition speed (up to 41 times improvement) without sacrificing accuracy. Our proposed approach offers a highly promising solution for ultra-fast, real-time, and high-precision dense depth estimation. Code and dataset will be publicly available.
Keywords:
Structured light Depth estimation Event camera Dynamic vision sensor
@@ 1 Introduction
Depth estimation plays a pivotal role in various fields[17, 42, 7], including machine vision and robotics, with applications spanning 3D modeling, autonomous driving, industrial inspection, and more. As a depth estimation method based on the principle of triangulation, the structured light system (SL system) projects a known pattern onto the target and estimates depth based on the pattern spatially-modulated by the target[2, 36]. The SL system is favored for its high accuracy, dense depth estimation capability, insensitivity to target texture features, and reasonable cost, rendering it indispensable in domains such as 3D scanning and augmented reality[45]. Presently, the SL system boasts numerous mature commercial applications, underscoring its widespread adoption and effectiveness.
@@
However, SL systems employing traditional frame-based cameras are restricted to low frequency (usually at a frame rate of tens of hertz) due to bandwidth and exposure time constraints of the cameras. While high-speed cameras offer the potential for ultra-fast acquisition[21], their exorbitant cost renders them impractical for many applications. Moreover, traditional SL systems suffer from data redundancy. Images captured by frame-based cameras contain substantial redundant information, particularly in point and line-structured light systems [1]. This redundancy undermines the acquisition speed and overall efficiency of the system.
@@
In recent years, event cameras[3, 10, 24], also known as Dynamic Vision Sensors (DVS), have emerged as a promising option for SL systems. These bio-inspired devices asynchronously report changes in light intensity at the $\mu$s level. Due to their capability of solely generating signals for changes in light intensity, event cameras inherently mitigate data redundancy in most SL systems at the hardware level. Additionally, the high dynamic range (HDR) of event cameras allows them to operate effectively under diverse illumination conditions. Consequently, event cameras offer an alternative solution for high-speed depth estimation in SL systems.
@@
Some works have already explored its potential in depth estimation. Laser-point projector based approaches[26, 30, 28] have achieved a scanning speed of 60Hz and high recovery accuracy. After program engineering, they can perform real-time dense target depth estimation at lower frame rates. However, this method is limited by the low encoding efficiency and timestamp noise (Latency and Jitter) , making it hard to achieve higher-speed depth estimation. On the other hand, works based on the full-frame method[23, 25, 40] can achieve higher speed by using Digital Light Processing (DLP) projector. However, these methods are still affected by timestamp noise and face difficulties in robust decoding.
@@
Therefore, considering the high-speed response and low redundancy characteristics of event cameras, as well as the high-speed projection characteristics of DLP projectors, we propose our Structured light system based on Gray code with an event camera (SGE) , which uses Gray code to encode the target, and estimates its depth based on epipolar geometry. This method first utilizes spatio-temporal encoding instead of timestamp matching for point matching, allowing for faster DLP projection and not being affected by noise of event camera and projector. Secondly, the binary nature of Gray code perfectly match with the binary response of event cameras. This allows our method to preserve all encoding information without data redundancy. Thirdly, we have developed a novel calibration scheme for event-based SL systems. Compared to previous schemes[27, 39, 9], our scheme has sub-pixel level accuracy and is convenient to use, and it also applicable to all event-based SL systems. Finally, we also propose the Gray code X-map (GX-map) disparity query scheme. This scheme enable us to directly query disparities instead of disparity searching. We achieved a processing speed improvement of three orders of magnitude in our experiment. These advantages allow us to achieve a much higher depth estimation speed exceeding 1000 Hz while maintaining comparable accuracy to state-of-the-art methods at a millimeter-level precision, and within a range of 1/200 of the average depth.
@@
  • Structured light system based on Gray code with an event camera system (SGE) is proposed with depth acquisition speed over 1000 Hz (Up to 41 times improvement). Experiments on both static and high-speed dynamic scenes show the superiority of the proposed SGE.
  • A sub-pixel level, universally applicable calibration method and the GX-map disparity query scheme are proposed. The latter improves the processing speed by a factor of 1000.
[…]
@@ 2 Related work
Structured light system. Traditional SL system performs depth estimation based on patterns spatial modulated by the target. As a field that has been researched for many years, it has yielded many accomplishments. Such as different projection pattern: point[1, 6, 14], line[5], patterned area[29, 49], binary coding[19, 16, 44], continuous coding[22, 15]. And as well as different numbers of pattern, including multiple-shot[47] and single-shot[38, 20, 48]. These schemes have their own advantages and disadvantages in terms of accuracy, speed, environmental adaptability, etc., thus requiring selection based on specific circumstances[13, 46]. However, due to the limited frame rate of frame-based cameras, typically only tens of Hz, it is impossible for traditional SL systems to achieve high-speed depth estimation. This limitation prevents their application in dynamic scenes.
@@
Event-based depth estimation. Due to advantages[10] of event cameras, for example high temporal resolution, they have been gradually used in the field of depth estimation. Initial approaches to event-based depth estimation primarily relied on feature matching with multi-view stereo techniques [32]. Subsequently, event-based Simultaneous Localization and Mapping (SLAM) methods [33, 35, 50] gained traction for constructing environmental representations or maps and deriving metric depth information. However, these approaches rely on either the scale provided by available camera poses or auxiliary sensors like inertial measurement units (IMUs) for metric depth recovery. Gallego et al. [11] removed dependencies on other modalities with a contrast maximization framework. Later, learning-based methods utilized events alone to predict depth with multi-view and egomotion constraints. Ye et al. [43] proposed a self-supervised framework for depth, optical flow, and egomotion prediction based on the principle of the egomotion model, which was further improved by Zhu et al. [51] with Temporal Reprojection Loss, Stereo Disparity Loss, and other constraints. These methods all utilized multi-view approaches, either stereo cameras or egomotion. Hidalgo et al. [18] were the first to propose predicting dense depth with a monocular camera using supervised learning. Gehrig et al. [12] combined frames and events for more robust depth estimation. However, learning-based methods still have significant limitations in terms of the accuracy of depth estimation. This heavily rely on high-quality datasets and proper training techniques. Furthermore, since the data is generated from slow target motion and need to pass through a slow network, both the data collection and processing speed are relatively slow.
@@
Event-based SL system. Event cameras have also gained attention in SL systems. Early work[4] used a 500Hz pulsed laser to achieve three-dimensional reconstruction by moving the target. However, this method required a displacement stage for scanning. Matsuda etal.[26] proposed Motion Contrast 3D (MC3D) replacing the pulsed laser with a laser-point projector and utilizing pixel timestamps for point matching. This approach laid a foundation for subsequent research but suffered from quality degradation under high-speed scanning due to timestamp noise. Muglikar etal.[30] analyzed this noise and improved MC3D’s matching algorithm, enhancing robustness against timestamp noise at a scanning speed of 60Hz. However, it slowed down processing speed due to increased computational effort. Morgenstern etal.[28] introduced X-map method, which used table look-up operations for fast disparity calculation and achieved real-time depth estimation. However, timestamp noise interference remained a challenge. Other works [23, 25, 40] explored DLP projection for faster depth estimation but faced issues of realizing high accuracy and robust decoding. In summary, most existing works struggle to achieve high-accuracy, high-speed, and robust depth estimation simultaneously, but showing favorable results only in specific aspects.
@@ 3 Method
We first introduce the foundation and basic consideration of our SGE method in Sec. 3.1 and then outline the pipeline in Sec. 3.2. In Sec. 3.3, our simple and high accuracy calibration scheme for event-based SL system will be proposed. Finally, we detail the depth estimation scheme based on GX-map query and Time-overlapping strategy in Sec. 3.4.
@@ 3.1 Preliminaries and Analysis
SL system for depth estimation. Traditional SL systems, composed of a frame-based camera and a projector, typically employ triangulation for depth estimation. Similar to common stereo vision system, SL systems can be conceptualized as artificially projecting "features" onto the target surface using a projector. Due to the typically dense nature of this projection, SL systems can achieve dense depth estimation that surpasses the capabilities of stereo vision alone, particularly in generating highly accurate depth information at short distances [37]. The primary distinction among SL schemes lies in the use of different projection patterns. But these patterns only lead to variations in depth encoding and do not impact the disparity calculation process. Therefore, in structured light depth estimation, after obtaining the depth encoding, stereo rectification is first performed based on the calibration parameters, and then point matching:
@@
where $P_{cr}$ denotes an encoded point in the camera’s pixel coordinate system, $P_{pr}$ represents the corresponding point in the projector’s pixel coordinate system. $x_{pr}$, $y_{pr}$, $x_{cr}$, $y_{cr}$ represent pixel coordinates in the corresponding pixel coordinate system. And $F$ represents the matching algorithm. Assuming that the disparity direction is along the $X$ direction, the depth can be obtained by:
@@
Here, $f$ denotes the focal length, while $b$ represents the baseline distance. However, SL systems encounter two basic limitations preventing it from high-speed measurement. First, it is restricted by low frame rate of the frame-based camera, typically only tens of hertz. Second, the abundant data redundancy of images not only takes up sensor bandwidth but also demands additional data processing[30].
@@
$p_{e,\tau}=\begin{cases}+1,\text{if}\log\left(\frac{\mathcal{I}_{t,e}}{\mathcal{I}_{(t-\Delta t),e}}\right)>C_{positive},\\ -1,\text{if}\log\left(\frac{\mathcal{I}_{t,e}}{\mathcal{I}_{(t-\Delta t),e}}\right)<C_{negative},\end{cases}\tau=t+\sigma$ (3)
where $\mathcal{I}$ represents intensity, $C_{positive}$, $C_{positive}$ represent the threshold for generating positive and negative events, which can be manually adjusted. $t$, $\sigma$, and $\tau$ are the ideal timestamp, the timestamp noise, and the actual output timestamp, respectively.
The previous works[30, 28] have utilized event-based SL system to realize depth estimation. However, these point scanning approaches suffer from timestamp noise and low encoding efficiency. The former leads to a decrease in accuracy when scanning speed increases. The latter means more timestamps are needed for depth encoding, as shown in Equ. 4, where N is the number of timestamps needed for encoding, C is the number of columns (disparity direction).
$N=C,\vspace{-1.5em}$ (4)
@@
Therefore, for the first time we propose using Gray code as depth encoding pattern for depth estimation in event-based SL systems. Firstly, by using Gray code, we achieve higher encoding efficiency while maintaining the absence of redundant information, as shown in Equ. 5. Secondly, our approach is immune to timestamp noise because we only need timestamps for data segmentation. Therefore, our SGE has the capacity to higher-speed and high-precision depth estimation.
Figure 1: The pipeline of the proposed SGE. The overall process starts with system calibration, and obtains depth after data acquisition and data processing.
@@ 3.2 Overall Pipeline
Our SGE comprises a DLP projector and an event camera, as depicted in Fig. 1. Before deployment, we first calibrate the system (Sec. 3.3) using our proposed Simple Event-based SL system Calibration (SEC). Then the projector independently and continuously projects Gray code onto the target. Thanks to DLP’s expertise in binary projection, our system can achieve an exceptionally rapid projection speed of up to 2487HZ, which is only limited by our hardware. Patterns modulated by the target depth are captured by the event camera from another view. As the data from the event camera is in the form of an event stream, segmentation according to each projection and subsequent conversion into image form are required (Step (a)). After acquiring a whole set of patterns, we use the traditional Gray code algorithms[31] to derive the depth-encoded image (Step (b)). Finally, leveraging the system’s calibrated parameters (Step (c)) and GX-map (Step (d)) , we determine the depth in Sec. 4. For dynamic scenes, we propose the Time-overlapping strategy to achieve higher data utilization and depth estimation speed (Sec. 5). Alg. 1 provides the pseudo-code of the proposed SGE.
Figure 2: Pipeline of calibration for event-based SL system. Event data from different angle are needed for calibration.

### v2

@@
SGE: Structured Light System Based on Gray Code with an Event Camera
[…]
journal: opticajournalarticletype: Research Article
{abstract*}
Fast and accurate depth sensing has long been a significant research challenge. Event camera, as a device that quickly responds to intensity changes, provides a new solution for structured light (SL) systems. In this paper, we introduce Gray code into event-based SL systems for the first time. Our setup includes an event camera and a Digital Light Processing (DLP) projector, enabling depth estimation through high-speed projection and decoding of Gray code patterns. By employing Gray code for point matching in event-based SL system, our method is immune to timestamp noise, realizing high-speed depth estimation without loss of accuracy and spatial resolution. The binary nature of events and Gray code minimizes data redundancy, enabling us to fully utilize sensor bandwidth at 100%. Experimental results show that our approach achieves accuracy comparable to state-of-the-art scanning methods while surpassing them in data acquisition speed (up to 41 times improvement) without sacrificing accuracy and spatial resolution. Our proposed approach offers a highly promising solution for ultra-fast, real-time, and high-precision dense depth estimation.
@@ 1 Introduction
Depth estimation plays a pivotal role in various fields[1, 2, 3], including machine vision and robotics, with applications spanning 3D modeling, autonomous driving, industrial inspection, and more. As a depth estimation method based on the principle of triangulation, the structured light system (SL system) projects a known pattern onto the target and estimates depth based on the pattern spatially modulated by the target[4, 5]. The SL system is favored for its high accuracy, dense depth estimation capability, insensitivity to target texture features, and reasonable cost, rendering it indispensable in domains such as 3D scanning and augmented reality[6]. Presently, the SL system boasts numerous mature commercial applications, underscoring its widespread adoption and effectiveness.
@@
However, SL systems employing traditional frame-based cameras are restricted to low frequency (usually at a frame rate of tens of hertz) due to bandwidth and exposure time constraints of the cameras. While high-speed cameras offer the potential for ultra-fast acquisition[7, 8, 9], their exorbitant cost renders them impractical for many applications. Moreover, traditional SL systems suffer from data redundancy. Images captured by frame-based cameras contain substantial redundant information, particularly in point and line-structured light systems [10]. This redundancy undermines the acquisition speed and overall efficiency of the system.
@@
Recently, event cameras[11, 12, 13, 14, 15, 16, 17, 18], also known as Dynamic Vision Sensors (DVS), have emerged as a promising option for SL systems. These bio-inspired devices asynchronously report changes in light intensity at the $\mu s$ level. Due to their capability of solely generating signals for changes in light intensity, event cameras inherently mitigate data redundancy in most SL systems at the hardware level. Additionally, the high dynamic range (HDR) of event cameras allows them to operate effectively under diverse illumination conditions. Consequently, event cameras offer an alternative solution for high-speed depth estimation in SL systems.
@@
Several studies have already explored its potential for depth estimation. Laser-point-projector-based approaches[19, 20, 21, 22] have achieved a scanning speed of 60 Hz and high recovery accuracy. After program engineering, they can perform real-time dense target depth estimation at lower frame rates. However, these methods are limited by the low encoding efficiency, timestamp noise (Latency and Jitter) and low scanning rate of the projector, making them hard to achieve higher-speed depth estimation. On the other hand, works based on the full-frame method can achieve higher speed by using Digital Light Processing (DLP) projectors. Among these methods, current temporal coding approaches [23, 24, 25] are still influenced by timestamp noise and face challenges in robust temporal decoding. Conversely, the spatial coding method [26] is a one-shot approach that achieves good results in the planar scene lacking detail, but at the cost of the resolution. The trade-off makes it difficult to capture detailed information and results in edge flickering due to spatial decoding.
@@
Therefore, considering the high-speed response and low redundancy characteristics of event cameras, as well as the high-speed projection characteristics of DLP projectors, we propose our Structured light system based on Gray code with an event camera (SGE), a temporal coding method which only uses Gray code to encode the target, and estimates its depth based on epipolar geometry. This method first utilizes temporal coding of Gray code, instead of using timestamp matching for point matching or needing accurate timestamps to get a phase map. In contrast to timestamp matching, our SGE enables faster DLP projection and is not influenced by the noise of the event camera and projector. In comparison with phase map methods (current temporal coding methods), our SGE offers robust temporal decoding, faster processing speed, and is also not affected by noise. Secondly, the binary nature of Gray code perfectly corresponds to the binary response of event cameras. This allows our method to preserve all encoding information without data redundancy. Thirdly, we have developed a novel calibration scheme for event-based SL systems. Compared to previous schemes[27, 28, 29], our scheme has sub-pixel level accuracy and is convenient to use, and it is also applicable to all event-based SL systems. Finally, we also propose the Gray code X-map (GX-map) disparity query scheme. This scheme enables us to directly query disparities instead of disparity searching. We achieved a processing speed improvement of around 250 times in our experiment. These advantages allow us to achieve a much higher depth estimation speed exceeding 1000 Hz while maintaining comparable accuracy to state-of-the-art methods at a millimeter-level precision, which is within a range of 1/200 of the average depth. Meanwhile, the straightforward yet effective pipeline enables our SGE to achieve real-time processing at approximately 200 Hz in 720 * 1280 resolution.
@@
  • A structured light system based on Gray code with an event camera system (SGE) is proposed with a depth acquisition speed over 1000 Hz (Up to 41 times improvement). Experiments on both static and high-speed dynamic scenes show the superiority of the proposed SGE.
  • A sub-pixel level, universally applicable calibration method, and the GX-map disparity query scheme are proposed. The latter improves the processing speed by a factor of around 250.
  • An optimized depth estimation pipeline has been put forward, enabling an algorithm capable of real-time processing.
@@ 2 Related work
Structured light system. Traditional SL system performs depth estimation based on patterns spatial modulated by the target. As a field that has been researched for many years, it has yielded many accomplishments. Such as different projection pattern: point[10, 30, 31], line[32], patterned area[33, 34], binary coding[35, 36, 37], continuous coding[38, 39]. And as well as different numbers of patterns, including multiple-shot[40] and single-shot[41, 42, 43]. These schemes have their advantages and disadvantages in terms of accuracy, speed, environmental adaptability, etc., thus requiring selection based on specific circumstances[44, 45]. However, due to the limited frame rate of frame-based cameras, typically only tens of Hz, it is impossible for traditional SL systems to achieve high-speed depth estimation. This limitation prevents their application in dynamic scenes.
@@
Event-based depth estimation. Due to advantages[12] of event cameras, for example, high temporal resolution, they have been gradually used in the field of depth estimation. Initial approaches to event-based depth estimation primarily relied on feature matching with multi-view stereo techniques [46]. Subsequently, event-based Simultaneous Localization and Mapping (SLAM) methods [47, 48, 49] gained traction for constructing environmental representations or maps and deriving metric depth information. However, these approaches rely on either the scale provided by available camera poses or auxiliary sensors like inertial measurement units (IMUs) for metric depth recovery. Gallego et al. [50] removed dependencies on other modalities with a contrast maximization framework. Later, learning-based methods utilized events alone to predict depth with multi-view and egomotion constraints. Ye et al. [51] proposed a self-supervised framework for depth, optical flow, and egomotion prediction based on the principle of the egomotion model, which was further improved by Zhu et al. [52] with Temporal Reprojection Loss, Stereo Disparity Loss, and other constraints. These methods all utilized multi-view approaches, either stereo cameras or egomotion. Hidalgo et al. [53] were the first to propose predicting dense depth with a monocular camera using supervised learning. Gehrig et al. [54] combined frames and events for more robust depth estimation. However, learning-based methods still have significant limitations in terms of the accuracy of depth estimation. This heavily rely on high-quality datasets and proper training techniques. Furthermore, since the data is generated from slow target motion and need to pass through a slow network, both the data collection and processing speed are relatively slow.
@@
Event-based SL system. Event cameras have also gained attention in SL systems. Early work[55] used a 500 Hz pulsed laser to achieve three-dimensional reconstruction by moving the target. However, this method required a displacement stage for scanning. Matsuda et al. [19] proposed Motion Contrast 3D (MC3D) replacing the pulsed laser with a laser-point projector and utilizing pixel timestamps for point matching. This approach laid a foundation for subsequent research but suffered from quality degradation under high-speed scanning due to timestamp noise. Muglikar et al. [20] analyzed this noise and improved MC3D’s matching algorithm, enhancing robustness against timestamp noise at a scanning speed of 60 Hz. However, it slowed down processing speed due to increased computational effort. Morgenstern et al. [21] introduced the X-map method, which used table look-up operations for fast disparity calculation and achieved real-time depth estimation. However, timestamp noise interference remained a challenge. Other works explored DLP projection for faster depth estimation. Huang et al. [26] employs a spatial coding method by using pseudo-random black-and-white dots and realizes a scan rate under 1000 Hz. On the measurement of the simple planar scene, they use the plane fitted by depth points they obtained as ground truth and calculate the root mean square error (RMSE) of measurement points and fitting points as their accuracy metrics, and achieve good results. However, this method further reduces the resolution when the resolution of the event camera is already not high, which makes it difficult to capture detailed information. It also suffers from edge flickering and complex data processing resulting from spatial decoding. Current works[23, 24, 25] based on temporal coding method faced issues of realizing high accuracy and robust decoding. In summary, most existing works struggle to achieve high accuracy, high speed, and robust depth estimation simultaneously but show favorable results only in specific aspects.
@@ 3 Method
We first introduce the foundation and basic consideration of our SGE method in Sec. 3.1 and then outline the pipeline in Sec. 3.2. In Sec. 3.3, our simple and high-accuracy calibration scheme for event-based SL systems will be proposed. Finally, we detail the depth estimation scheme based on the GX-map query and Time-overlapping strategy in Sec. 3.4.
@@ 3.1 Preliminaries and Analysis
SL system for depth estimation. Traditional SL systems, composed of a frame-based camera and a projector, typically employ triangulation for depth estimation. Similar to common stereo vision systems, SL systems can be conceptualized as artificially projecting "features" onto the target surface using a projector. Due to the typically dense nature of this projection, SL systems can achieve denser depth estimation, surpassing stereo vision, especially in generating highly accurate depth information at short distances[56]. The primary distinction among SL schemes lies in the use of different projection patterns. However, these patterns only lead to variations in depth encoding and do not impact the disparity calculation process. Therefore, in structured light depth estimation, after obtaining the depth encoding, stereo rectification is first performed based on the calibration parameters, and then point matching:
@@
where $P_{cr}$ denotes an encoded point in the camera’s pixel coordinate system, and $P_{pr}$ represents the corresponding point in the projector’s pixel coordinate system. $x_{pr}$, $y_{pr}$, $x_{cr}$, $y_{cr}$ represent pixel coordinates in the corresponding pixel coordinate system. And $F$ represents the matching algorithm. Assuming that the disparity direction is along the $X$ direction, the depth can be obtained by:
@@
Here, $f$ denotes the focal length, while $b$ represents the baseline distance. However, SL systems encounter two basic limitations preventing them from high-speed measurement. First, it is restricted by low frame rate of the frame-based camera, typically only tens of hertz. Second, the abundant data redundancy of images not only takes up sensor bandwidth but also demands additional data processing[20].
Gray code SL system. The Gray code SL system uses patterns with binary stripes for projection[44]. Thus, each point along the disparity direction has a unique binary code for later point matching. Its binary nature makes the method robust and adaptable. However, this nature also leads to low grayscale information utilization compared to other grayscale-based methods, requiring a large number of patterns for high spatial resolution, which increases data acquisition time. But with an event camera, only binary information is obtained from the hardware to speed up data acquisition, making Gray code perfect for the event-based system. All the above answers the question of why this work combines the Gray code and the event camera.
@@
$p_{e,\tau}=\begin{cases}+1,\text{if}\log\left(\frac{\mathcal{I}_{t,e}}{\mathcal{I}_{(t-\Delta t),e}}\right)>C_{positive},\\ -1,\text{if}\log\left(\frac{\mathcal{I}_{t,e}}{\mathcal{I}_{(t-\Delta t),e}}\right)<C_{negative},\end{cases}$ (3)
where
$\tau=t+\sigma,$ (4)
$\mathcal{I}$ represents intensity, $C_{positive}$ and $C_{negative}$ represent the threshold for generating positive and negative events, which can be manually adjusted. $t$, $\sigma$, and $\tau$ are the ideal timestamp, the timestamp noise, and the actual output timestamp, respectively.
The previous works[20, 21, 26, 23, 24, 25] have utilized event-based SL systems to realize depth estimation. However, point-scanning approaches suffer from timestamp noise and low encoding efficiency. The former leads to a decrease in accuracy when scanning speed increases. The latter means more timestamps are needed for depth encoding, as shown in Equ. 5, where $n$ is the number of timestamps needed for encoding, and $c$ is the number of columns (disparity direction). The temporal coding has issues such as not robust temporal decoding and is still affected by timestamp noise. Meanwhile, the spatial coding has other problems like low spatial resolution and edge flickering.
$n=c,$ (5)
@@
Therefore, for the first time, we propose to use Gray code as a depth encoding pattern for depth estimation in event-based SL systems. Firstly, compared to point scanning methods, by using Gray code, we achieve robust temporal decoding and higher encoding efficiency while maintaining the absence of redundant information, as shown in Equ. 6. Secondly, our approach is immune to timestamp noise because we only need timestamps for data segmentation. Thirdly, compared with the spatial coding methods, which sacrifice spatial resolution for acquisition speed, our approach remains the origin spatial resolution. These illustrate the capacity of our SGE to achieve high speed, high-precision depth estimation, and high-resolution details.
@@ 3.2 Overall Pipeline
Figure 1: The pipeline of the proposed SGE. Patterns are loaded and cast onto a target, and then captured by an event camera in the event stream type continuously. Then binarized images are obtained by conducting step (5). When a set of images is gathered, steps (6-8) are conducted to estimate depth.
Our SGE comprises a DLP projector and an event camera, as depicted in Fig. 1. Before deployment, we first calibrate the system (Sec. 3.3) using our proposed Simple Event-based SL system Calibration (SEC). Then the projector independently and continuously projects Gray code onto the target. Thanks to DLP’s expertise in binary projection, our system can achieve an exceptionally rapid projection speed of up to 2487 Hz, which is only limited by our hardware. Patterns modulated by the target depth are captured by the event camera from another view. As the data from the event camera is in the form of an event stream, segmentation according to each projection and subsequent conversion into image form are required. After acquiring a whole set of patterns, we use the traditional Gray code algorithms[44, 57] to derive the depth-encoded image. Finally, leveraging the system’s calibrated parameters and GX-map, we determine the depth in Sec. 6. For dynamic scenes, we propose the Time-overlapping strategy to achieve higher data utilization and depth estimation speed (Sec. 6). We also conduct experiments on the low computational cost of our proposed SGE in Sec. 4.5 to show the real-time processing capability. Finally, Alg. 1 shows the pseudo-code of the proposed SGE.
Figure 2: The calibration pipeline of our SEC. The event camera captures a projected checkerboard and a flickering checkerboard on the same plane from various perspectives. Then binarizing them into images and using the Gabriel calibration tool to calibrate SL systems.